# In search of maximal branes on hyper-Kähler manifolds

### v1

@@ Abstract
Given a holomorphic or anti-holomorphic involution on a complex variety, the Smith inequality says that the total $\operatorname{\mathbb{F}_{2}}2$-Betti number of the fixed locus is no greater than the total $\operatorname{\mathbb{F}_{2}}2$-Betti number of the ambient variety. The involution is called maximal when the equality is achieved. In this paper, we investigate the existence problem for maximal involutions on higher-dimensional compact hyper-Kähler manifolds and on Hilbert schemes of points on surfaces.
We show that for $n\geq 2$, a hyper-Kähler manifold of K3[n]-deformation type admits neither maximal anti-holomoprhic involutions (i.e. real structures), nor maximal holomorphic (symplectic or anti-symplectic) involutions. In other words, such hyper-Kähler manifolds do not contain maximal (AAB), (ABA), (BAA) or (BBB)-branes.
For Hilbert schemes of points on surfaces, we show that for a holomorphic (resp. anti-holomorphic) involution $\sigma$ on a smooth projective surface $S$ with $H^{1}(S,\operatorname{\mathbb{F}_{2}}2)=0$, the naturally induced involution on the $n$th Hilbert scheme of points is maximal if and only if $\sigma$ is a maximal involution of $S$ and it acts on $H^{2}(S,\mathbb{Z})$ trivially (resp. as $\operatorname{id}$). This generalizes previous results of Fu and Kharlamov–Răsdeaconu.
@@ 1.2. Motivation from real geometry
When equality in (1.2) holds, we call $(X,\sigma)$ a maximal real variety (or M-variety). Maximal real varieties have attracted significant research interest over the decades. Let us mention the remarkable Rokhlin congruence theorem for even-dimensional maximal smooth projective real varieties: $\chi(X(\mathbb{R}))\equiv\operatorname{sgn}(X)\mod 16$; see for example [38, Theorem 3.4.2]. On the other hand, constructing examples of maximal real varieties in dimension >2>2 is often challenging; see [5, §3] and [17] for a recent summary and see [38] for the case of curves and surfaces. In higher dimensions, Viro’s combinatorial patchworking is very powerful for complete intersections [29]. Let us mention the second author’s recent contribution [17] of constructing new maximal real varieties by looking at moduli spaces. The present paper is a natural continuation of [17], with a more focused investigation on hyper-Kähler manifolds and Hilbert schemes  [^2] It should rather be called Douady space in the complex analytic category. of points on surfaces.
1.3. Branes in hyper-Kähler manifolds
@@
On a hyper-Kähler manifold $X$, a real structure, or equivalently an anti-holomorphic involution $\sigma$, is referred to as an (ABA) or (AAB) brane involution [^3] The distinction between (ABA) and (AAB) depends on the action of $\sigma$ on a choice of a (unique up to scalar) holomorphic symplectic form $\eta$: it is called (ABA) if $\sigma^{*}(\eta)=\overline{\eta}$, and called (AAB) if $\sigma^{*}(\eta)=-\overline{\eta}$., and its fixed locus, which is precisely the real locus $X(\mathbb{R})$, is known as an (ABA) or (AAB) brane in $X$, as defined in [1] (see also [15, §2.3]). Up to hyper-Kähler rotation of the complex structure, an anti-holomorphic involution can become a holomorphic anti-symplectic involution, referred to as a (BAA)-brane involution. Similarly, a holomorphic symplectic involution on a hyper-Kähler manifold is referred to as a (BBB)-involution and its fixed locus is called a (BBB)-brane.
@@
Examples of non-compact hyper-Kähler manifolds with maximal (ABA) or (AAB) branes are constructed via moduli spaces of Higgs bundles by Fu in [17, Theorem 6.3]. Another example of maximal brane in a non-compact hyper-Kähler manifold is provided by the cotangent bundle of a maximal $\mathbb{R}$-variety.
@@
However, the situation is more intriguing for compact hyper-Kähler manifolds. We refer to [2] and [24] for generalities of such manifolds; see also Section˜4 for a quick summary of their special properties. Let us just mention here the most studied examples of compact hyper-Kähler manifolds: deformations of Hilbert schemes of points on K3 surfaces; such hyper-Kähler manifolds are called of $\operatorname{\mathrm{K3}}3^{[n]}$-type. On the one hand, in (complex) dimension 2, maximal real K3 surfaces and abelian surfaces exist and have been thoroughly studied; see [31], [48], [53, Chapters IV, VIII]. On the other hand, Kharlamov and Răsdeaconu [32] made the surprising discovery that the Hilbert square of maximal real K3 surfaces (see also [33] for related results) and Fano varieties of lines of maximal real cubic fourfolds are never maximal.
The motivation of this paper is to investigate the following question, raised by Fu in [17]:
@@
Can compact hyper-Kähler manifolds of dimension $\geq 4$ admit maximal (ABA), (AAB), (BAA) and (BBB) branes?
@@
1.4. Main results I: non-existence of maximal branes in hyper-Kähler manifolds of $\operatorname{\mathrm{K3}}3^{[n]}$-type
@@
Although there are K3 surfaces and abelian surfaces admitting maximal brane involutions, our first main result proves the non-existence of maximal (ABA), (AAB) or (BAA) branes in the compact hyper-Kähler manifolds of $\operatorname{\mathrm{K3}}3^{[n]}$-type.
Theorem 1.2 (Absence of maximal (BAA)-brane).
@@
By hyper-Kähler rotation, we immediately get the following application in real algebraic geometry, answering a question raised by the second author in [17] for these most studied deformation families of compact hyper-Kähler manifolds.
Corollary 1.3 (Absence of maximal (ABA)/(AAB)-brane).
@@
We also prove the parallel result for (BBB)-branes.
Theorem 1.4 (Absence of maximal (BBB) branes).
@@
Remark 1.5.
Theorem˜1.2, ˜1.3, Theorem˜1.4 provide convincing evidence on the non-existence of maximal (ABA), (AAB), (BAA), and (BBB) branes for compact hyper-Kähler manifolds of dimension larger than $2$. It would be extremely interesting to exploit various special properties of compact hyper-Kähler manifolds to give an a priori reason to the non-existence of such maximal branes.
@@ 1.5. Main results II: criteria for maximality of natural involutions on Hilbert schemes
To motivate our second main result, we extract here the following special case of Theorem˜1.2, ˜1.3 and Theorem˜1.4.
@@
Going beyond hyper-Kähler geometry, recall that the Hilbert scheme of points on a smooth surface is always smooth [14], and a holomorphic or anti-holomorphic involution on the surface naturally induces an involution on the Hilbert scheme of points by base-change (cf. [7, 5.3.1] for the anti-holomorphic case). We regard ˜1.6 as an example towards the investigation of the more general question whether the operations $^{[n]}$ of taking Hilbert powers on a surface preserve the maximality:
Question 1.7.
Given a smooth surface $S$ equipped with a holomorphic or anti-holomorphic involution, when is the natural involution on the Hilbert scheme of $n$ points $S^{[n]}$ maximal?
This question has been recently studied in Fu [17] and in Kharlamov–Răsdeaconu [32]. For projective surfaces satisfying $H^{1}(S,\operatorname{\mathbb{F}_{2}}2)=0$, our results below provide a complete answer to ˜1.7 by giving a sufficient and necessary condition, solely in terms of the involution on $S$.
@@
Theorem˜1.8 establishes the converse of [17, Theorem 8.1] and generalizes [32, Theorem 1.2] to all $n\geq 2$. In a similar fashion, the following theorem gives a clean characterization for the maximality of natural holomorphic involution on punctual Hilbert schemes of surfaces. In Section˜6.3 these theorems are applied to obtain examples of maximal involutions and non-maximal involutions, generalizing [32, Corollaries 1.3 and 1.4].
@@ 1.6. Outline of the article
The structure of the article is the following. Section˜2 contains preliminaries about maximal involutions, including sufficient and necessary conditions for maximality that are exploited throughout the paper. In Section˜3 we recollect facts about the basis given by Nakajima and Li–Qin–Wang for the integral cohomology of the Hilbert scheme of points on a smooth projective surface with vanishing $H^{1}(-,\operatorname{\mathbb{F}_{2}}2)$. Section˜4 is an introduction to hyper-Kähler manifolds with a particular focus on manifolds of K3[n]-type, their monodromy representation on the cohomology of degree 4, and the monodromy representation in terms of the integral basis for Hilbert schemes of points on K3 surfaces introduced in Section˜3. Section˜5 is dedicated to a proof of the fact that for a surface with $H^{1}(S,\mathbb{F}_{2})=0$ and a free holomorphic or anti-holomorphic involution, the induced involution on the Hilbert scheme of points is not maximal. The proof relies on tools of algebraic topology, such as the Smith–Gysin exact sequence and the Kalinin spectral sequence, together with the classification of complex surfaces. The content of Section˜6 is the proofs of Theorem˜1.8 and Theorem˜1.9. These results are achieved using the necessary and sufficient conditions of Section˜2 together with the facts about the integral basis for the cohomology of Hilbert schemes of points of Section˜3. Section˜7 contains the proof of the non-existence of maximal branes for hyper-Kähler manifolds of K3[n]-type, namely the proof of Theorem˜1.2, ˜1.3 and Theorem˜1.4. The main ingredients of the proofs are the necessary and sufficient conditions for maximality from Section˜2 and the monodromy representation on the degree 4 cohomology from Section˜4, together with some lattice theory. Some final comments and questions are collected in Section˜8.
@@ Acknowledgments
The $n=2$ case of Theorem˜1.2 was obtained earlier by the fourth author V. Kharlamov, who has benefited from discussions with Rares Răsdeaconu. We all thank Răsdeaconu for his contribution. We also thank Olivier Benoist for helpful conversations.
@@ 2. Preliminaries on maximality of involutions
Proposition 2.1 ([54, Chapter III, Proposition 4.16]).
@@
The following classification of integral representations of involutions is going back to Comessatti [9, 10] (for a sketch of proof see, for example, [53, Chapter I, Lemmas 3.5, 3.5.1]).
@@ 3. Integral cohomology of Hilbert schemes of points on surfaces
In the rest of the paper, cohomological correspondences will be frequently used. Let us recall the definition, which works for any choice of coefficients (cf. [45, Section 8.1]). Given two smooth projective varieties $X,Y$ and a cohomology class $\gamma\in H^{*}(X\times Y)$, we define
@@ 3.1. Nakajima operators and Li–Qin–Wang integral operators
Let $S$ be a smooth projective surface over $\mathbb{C}$. Following [51, Definition 3.1], an operator in $\operatorname{End}\left(\bigoplusop\displaylimits_{n}H^{*}(S^{[n]},\mathbb{Q})\right)$ is called integral, if it respects the subgroup $\bigoplusop\displaylimits_{n}H^{*}(S^{[n]},\mathbb{Z})/\mathrm{tors}\subset\bigoplusop\displaylimits_{n}H^{*}(S^{[n]},\mathbb{Q})$. This notion is slightly weaker than the usual one (elements in $\operatorname{End}(\bigoplusop\displaylimits_{n}H^{*}(S^{[n]},\mathbb{Z}))$).
@@
    For any $k\geq 0$, Li–Qin–Wang [36, Definition 4.1], [35, (2.7)] defined the operator $\mathds{1}_{-k}$ as $\mathds{1}_{-k}:=\frac{1}{k!}\mathfrak{p}_{-1}(1_{S})^{k}$. Despite of the apparent denominator, in [51, Lemma 3.3] it is proved that $\mathds{1}_{-k}$ is an integral operator. Indeed, it can be equivalently defined as follows: for any $j\geq 0$, the operator
@@
    The operator $\mathfrak{m}$ is defined in [51, §4.2] and in [35, Definition 3.3]. More precisely, for $\lambda$ a partition of $n$, $\alpha\in H^{2}(S,\mathbb{Z})/\mathrm{tors}$, we have an operator
@@
    which is proven to be an integral operator by Li–Qin in [35, Theorem 3.6]. The precise definition of $\mathfrak{m}$ is somewhat involved. Let us only mention here that when $\alpha=[C]$ for a smooth irreducible curve $C$ in $S$, then for any partition $\lambda=(\lambda_{1}\geq\cdots\geq\lambda_{N})$ of $n$, we have $\mathfrak{m}_{\lambda,C}|0\rangle=[L^{\lambda}C]$, where $L^{\lambda}C$ is the closure in $S^{[n]}$ of $\{\lambda_{1}x_{1}+\cdots+\lambda_{N}x_{N}~|~x_{i}\in C\text{ distinct}\}.$

### v2

# On maximality of involutions of hyper-Kähler manifolds and punctual Hilbert schemes of surfaces

@@ Abstract
Given a holomorphic or anti-holomorphic involution on a complex variety, the Smith inequality says that the total $\operatorname{\mathbb{F}_{2}}2$-Betti number of the fixed locus is no greater than the total $\operatorname{\mathbb{F}_{2}}2$-Betti number of the ambient variety. The involution is called maximal when the equality is achieved. In this paper, we investigate maximality of involutions of compact hyper-Kähler manifolds and of Hilbert schemes of points on surfaces. We obtain both positive and negative results.
On one hand, given a smooth projective surface $S$ with $H^{1}(S,\operatorname{\mathbb{F}_{2}}2)=0$ equipped with a holomorphic (resp. anti-holomorphic) involution $\sigma$, we establish the following necessary and sufficient condition for the maximality of the induced involution on the $n$th Hilbert scheme of points: the induced involution is maximal if and only if $\sigma$ is a maximal involution of $S$ and it acts on $H^{2}(S,\mathbb{Z})$ trivially (resp. as $\operatorname{id}$). This generalizes and completes previous partial results of Fu and Kharlamov–Răsdeaconu.
On the other hand, we show that for $n\geq 2$, a hyper-Kähler manifold of K3[n]-deformation type admits neither maximal anti-holomorphic involutions (i.e. real structures), nor maximal holomorphic (symplectic or anti-symplectic) involutions. In other words, such hyper-Kähler manifolds do not admit maximal (AAB), (ABA), (BAA) or (BBB) brane involutions in the sense of Kapustin–Witten.
@@ 1.2. Motivation from real geometry
When the equality in (1.2) holds, we call $(X,\sigma)$ a maximal real variety (or M-variety). Since the solution of the concrete questions on plane sextic curves and space quartic surfaces raised by D. Hilbert in his 16-th problem, achieved in the 1970’s ([26], [37]), maximal real varieties have entered, and remain, in the center of attention in real algebraic geometry. Many remarkable properties of these varieties were discovered, like Rokhlin’s congruence theorem for even-dimensional maximal smooth projective real varieties: $\chi(X(\mathbb{R}))\equiv\operatorname{sgn}(X)\mod 16$; see for example [44, Theorem 3.4.2]. However, constructing examples of maximal real varieties in dimension >2>2 still remains challenging; see [6, §3] and [20] for a recent summary and see [44] for the case of curves and surfaces. Even for hypersurfaces in projective spaces, despite the powerful Viro’s method of patchworking ([65], [66], [33]), it is still unknown whether maximal hypersurfaces exist for any degree in any dimension. Apart from hypersurfaces and complete intersections in projective spaces, another important class of varieties are various moduli spaces. Remarkably, they all have a natural real structure and their possible maximality is an important source of applications. That is the reason why in the last decades maximality of moduli spaces became a subject of research. Let us mention here the recent contributions by Etingof–Henriques–Kamnitzer–Rains [16], Brugallé–Schaffhauser [6] and Fu [20]. The present paper is a natural continuation of [20], with a more focused investigation on Hilbert schemes of points on surfaces and hyper-Kähler manifolds.
As is shown in the seminal works of Göttsche [23] and Nakajima [52], Hilbert schemes [^2] It should rather be called Douady space in the complex analytic category. of points on surfaces play distinctive roles among moduli spaces, in that this construction provides strong relations between the geometry of the surface and the geometry of an infinite series of varieties of arbitrarily high dimensions. Our initial motivation stems from the desire to determine when the Hilbert schemes of a real surface are maximal. More precisely, let us formulate the question in a boarder context as follows. Recall that the Hilbert scheme of points on a smooth surface is always smooth [17], and a holomorphic or anti-holomorphic involution on the surface naturally induces an involution on the Hilbert scheme of points by base-change (cf. [8, 5.3.1] for the anti-holomorphic case).
Question 1.2.
Given a smooth complex surface $S$ equipped with a holomorphic or anti-holomorphic involution, when is the naturally induced involution on the Hilbert scheme of $n$ points $S^{[n]}$ maximal?
Partial results towards this question have been recently obtained in Fu [20] and in Kharlamov–Răsdeaconu [39]. In this paper, we give a complete answer to it for projective surfaces satisfying $H^{1}(S,\operatorname{\mathbb{F}_{2}}2)=0$; see Section˜1.5.
1.3. Motivation from hyper-Kähler geometry: brane involutions
Given a Kähler manifold $X$ with complex structure $I$ and the associated symplectic form $\omega_{I}$, following Kapustin–Witten [36], an $A$-brane refers to a Lagrangian submanifold with respect to $\omega_{I}$, and a $B$-brane stands for a complex submanifold with respect to $I$.
For a hyper-Kähler manifold $(X,g)$ with three complex structures $I,J,K$ satisfying the quaternion relations, let $\omega_{I},\omega_{J},\omega_{K}$ be the associated Kähler forms. It is extremely interesting to study submanifolds of $X$ that are Lagrangian or complex with respect to each of the three Kähler structures. Such a submanifold is called an (AAB), (ABA), (BAA) or (BBB) brane, depending on its compatibility types with respect to the three Kähler structures. As is pointed out in [1] (see also [18, §2.3]), a natural and rich source of such branes in hyper-Kähler manifolds is provided by the fixed loci of so-called brane involutions — those are involutions preserving the hyper-Kähler metric such that each of the three complex structures is either preserved (B-type) or reversed (A-type).
@@ 1.3. Motivation from hyper-Kähler geometry: brane involutions
In concrete terms, on a hyper-Kähler manifold $X$, a real structure, or equivalently, an anti-holomorphic involution $\sigma$, gives rise to an (ABA) or (AAB) brane involution [^3] The distinction between (ABA) and (AAB) depends on the action of $\sigma$ on a choice of a (unique up to scalar) holomorphic symplectic form $\eta$: it is called (ABA) if $\sigma^{*}(\eta)=\overline{\eta}$, and called (AAB) if $\sigma^{*}(\eta)=-\overline{\eta}$., and its fixed locus, which is precisely the real locus $X(\mathbb{R})$, represents an (ABA) or (AAB) brane in $X$. Up to hyper-Kähler rotation of the complex structure, an anti-holomorphic involution can become a holomorphic anti-symplectic involution, i.e. a (BAA)-brane involution, and the fixed locus, which is a holomorphic Lagrangian submanifold, is a (BAA) brane. Similarly, a holomorphic symplectic involution on a hyper-Kähler manifold is referred to as a (BBB)-involution and its fixed locus represents a (BBB)-brane, also known as trianalytic submanifold introduced by Verbitsky [62].
@@
Examples of non-compact hyper-Kähler manifolds with maximal (ABA) or (AAB) brane involutions are constructed via moduli spaces of Higgs bundles by Fu in [20, Theorem 6.3]. Another example of maximal (ABA) or (AAB) brane involution in a non-compact hyper-Kähler manifold is provided by the natural real structure of the cotangent bundle of a maximal $\mathbb{R}$-variety.
@@
However, the situation is more intriguing for compact hyper-Kähler manifolds. We refer to [3] and [28] for generalities of such manifolds; see also Section˜4 for a quick summary of their special properties. Let us just mention here the most studied examples of compact hyper-Kähler manifolds: deformations of Hilbert schemes of points on K3 surfaces; such hyper-Kähler manifolds are called of $\operatorname{\mathrm{K3}}3^{[n]}$-type. On the one hand, in (complex) dimension 2, maximal real K3 surfaces and abelian surfaces exist and have been thoroughly studied; see [37], [54] [59, Chapters IV, VIII]. On the other hand, Kharlamov and Răsdeaconu [39] made the surprising discovery that the Hilbert square of maximal real K3 surfaces (see also [38] for related results) and Fano varieties of lines of maximal real cubic fourfolds are never maximal. These works lead us to the second motivation of this paper, summarized in the following question, raised by Fu in [20]:
Question 1.3.
@@
Can compact hyper-Kähler manifolds of dimension $\geq 4$ admit maximal (ABA), (AAB), (BAA) and (BBB) brane involutions?
In what follows, we give a decisive but rather unexpected answer to it for hyper-Kähler manifolds of $\operatorname{\mathrm{K3}}3^{[n]}$-type.
@@
1.4. Main results I: non-existence of maximal brane involutions on $\operatorname{\mathrm{K3}}3^{[n]}$-type manifolds
@@
Although there are K3 surfaces and abelian surfaces admitting maximal brane involutions, our first main result proves the non-existence of maximal (ABA), (AAB) or (BAA) brane involutions in the compact hyper-Kähler manifolds of $\operatorname{\mathrm{K3}}3^{[n]}$-type.
Theorem 1.4 (Absence of maximal (BAA)-brane involutions).
@@
By hyper-Kähler rotation, we immediately get the following application in real algebraic geometry, answering the aforementioned question raised by the second author in [20] for these most studied deformation families of compact hyper-Kähler manifolds.
Corollary 1.5 (Absence of maximal (ABA)/(AAB)-brane involutions).
@@
We also prove the parallel result for (BBB)-brane involutions.
Theorem 1.6 (Absence of maximal (BBB)-brane involutions).
@@ 1.5. Main results II: criteria for maximality of natural involutions on Hilbert schemes
Before giving our second main result, we extract here the following special case of Theorem˜1.4, ˜1.5 and Theorem˜1.6.
@@
Going beyond hyper-Kähler geometry, we regard ˜1.7 as an example towards the investigation of the more general ˜1.2 whether the operations $^{[n]}$ of taking Hilbert powers on a surface preserve the maximality. For projective surfaces $S$ with $H^{1}(S,\operatorname{\mathbb{F}_{2}}2)=0$, our results below provide a complete answer to ˜1.2 by giving a sufficient and necessary condition, solely in terms of the involution on $S$.
@@
Theorem˜1.8 establishes the converse of [20, Theorem 8.1] and generalizes [39, Theorem 1.2] to all $n\geq 2$. In a similar fashion, the following theorem gives a clean characterization for the maximality of natural holomorphic involution on punctual Hilbert schemes of surfaces. In Section˜6.3 these theorems are applied to obtain examples of maximal involutions and non-maximal involutions, generalizing [39, Corollaries 1.3 and 1.4].
@@ 1.6. Outline of the article
The structure of the article is the following. Section˜2 contains preliminaries about maximal involutions, including sufficient and necessary conditions for maximality that are exploited throughout the paper. In Section˜3 we recollect facts about the basis given by Nakajima and Li–Qin–Wang for the integral cohomology of the Hilbert scheme of points on a smooth projective surface with vanishing $H^{1}(-,\operatorname{\mathbb{F}_{2}}2)$. Section˜4 is an introduction to hyper-Kähler manifolds with a particular focus on manifolds of K3[n]-type, their monodromy representation on the cohomology of degree 4, and the monodromy representation in terms of the integral basis for Hilbert schemes of points on K3 surfaces introduced in Section˜3. Section˜5 is dedicated to a proof of the fact that for a surface with $H^{1}(S,\mathbb{F}_{2})=0$ and a free holomorphic or anti-holomorphic involution, the induced involution on the Hilbert scheme of points is not maximal. The proof relies on tools of algebraic topology, such as the Smith–Gysin exact sequence and the Kalinin spectral sequence, together with the classification of complex surfaces. The content of Section˜6 is the proofs of Theorem˜1.8 and Theorem˜1.9. These results are achieved using the necessary and sufficient conditions of Section˜2 together with the facts about the integral basis for the cohomology of Hilbert schemes of points of Section˜3. Section˜7 contains the proof of the non-existence of maximal brane involutions for hyper-Kähler manifolds of K3[n]-type, namely the proof of Theorem˜1.4, ˜1.5 and Theorem˜1.6. The main ingredients of the proofs are the necessary and sufficient conditions for maximality from Section˜2 and the monodromy representation on the degree 4 cohomology from Section˜4, together with some lattice theory. Some final comments and questions are collected in Section˜8.
@@ Acknowledgments
The $n=2$ case of Theorem˜1.4 was obtained earlier by the fourth author V. Kharlamov, who has benefited from discussions with Rares Răsdeaconu. We all thank Răsdeaconu for his contribution. We also thank Olivier Benoist for helpful conversations.
@@ 2. Preliminaries on maximality of involutions
Proposition 2.1 ([60, Chapter III, Proposition 4.16]).
@@
The following classification of integral representations of involutions is going back to Comessatti [10, 11] (for a sketch of proof see, for example, [59, Chapter I, Lemmas 3.5, 3.5.1]).
@@ 3. Integral cohomology of Hilbert schemes of points on surfaces
In the rest of the paper, cohomological correspondences will be frequently used. Let us recall the definition, which works for any choice of coefficients (cf. [52, Section 8.1]). Given two smooth projective varieties $X,Y$ and a cohomology class $\gamma\in H^{*}(X\times Y)$, we define
@@ 3.1. Nakajima operators and Li–Qin–Wang integral operators
Let $S$ be a smooth projective surface over $\mathbb{C}$. Following [57, Definition 3.1], an operator in $\operatorname{End}\left(\bigoplusop\displaylimits_{n}H^{*}(S^{[n]},\mathbb{Q})\right)$ is called integral, if it respects the subgroup $\bigoplusop\displaylimits_{n}H^{*}(S^{[n]},\mathbb{Z})/\mathrm{tors}\subset\bigoplusop\displaylimits_{n}H^{*}(S^{[n]},\mathbb{Q})$. This notion is slightly weaker than the usual one (elements in $\operatorname{End}(\bigoplusop\displaylimits_{n}H^{*}(S^{[n]},\mathbb{Z}))$).
@@
    For any $k\geq 0$, Li–Qin–Wang [42, Definition 4.1], [41, (2.7)] defined the operator $\mathds{1}_{-k}$ as $\mathds{1}_{-k}:=\frac{1}{k!}\mathfrak{p}_{-1}(1_{S})^{k}$. Despite of the apparent denominator, in [57, Lemma 3.3] it is proved that $\mathds{1}_{-k}$ is an integral operator. Indeed, it can be equivalently defined as follows: for any $j\geq 0$, the operator
@@
    The operator $\mathfrak{m}$ is defined in [57, §4.2] and in [41, Definition 3.3]. More precisely, for $\lambda$ a partition of $n$, $\alpha\in H^{2}(S,\mathbb{Z})/\mathrm{tors}$, we have an operator
@@
    which is proven to be an integral operator by Li–Qin in [41, Theorem 3.6]. The precise definition of $\mathfrak{m}$ is somewhat involved. Let us only mention here that when $\alpha=[C]$ for a smooth irreducible curve $C$ in $S$, then for any partition $\lambda=(\lambda_{1}\geq\cdots\geq\lambda_{N})$ of $n$, we have $\mathfrak{m}_{\lambda}([C])|0\rangle=[L^{\lambda}C]$, where $L^{\lambda}C$ is the closure in $S^{[n]}$ of $\{\lambda_{1}x_{1}+\cdots+\lambda_{N}x_{N}~|~x_{i}\in C\text{ distinct}\}.$